The number of subscribers to an internet streaming services is linearly related to the price. If the price is $10, there will be 4 million subscribers, but if the price is raised to $15 there will only be 3 million subscribers. Suppose the current price is $10, and the price is rising at $0.50 per month. Find the rate at which the revenue is changing at this time.
Answer #1

million -R-15 . D=10 at 4 = 4 - a :3 (4,10) (3,182 1o. 15-10 4 3-4 = -5(4-4 = -5x +30 ) P= Drama -5x+30 R: H. Dra) . 46-5x +dr.. (Clou+30) x du. Tdt dt = (-lout 3074 - 0. L giron Flox4 +30) Xiaol - 10x -0.1